# Identification of Semiparametric Panel Multinomial

### v2

@@ 2.2 Key Identification Strategy
Relative to the well-known maximum-score criterion function as studied by Manski (1985, 1987) utilizing univariate monotonicity, the nonstandardness of our criterion function arises from a key difference of multivariate monotonicity from univariate monotonicity. To see this more clearly, consider the special case of a single-index setting $(J=1)$ [^5] This arises naturally in binomial choice models with the characteristics of the outside option set to be zero. In this case, even though there are nominally two choice alternatives, choice behavior is completely determined by a single index based on the characteristics of the non-default option., in which case the following equivalence relationship holds given the univariate monotonicity in the index:
$\left\{\gamma\left(X_{t},X_{s}\right)>0\right\}\ \Leftrightarrow\ \left\{\left(X_{t}-X_{s}\right)\beta>0\right\},$ (12)
@@
Such an “if-and-only-if” relationship is a unique feature of the single-index setting, which cannot be generalized to the multi-index setting with $J\geq 2$, as the right hand side of (6),
@@
$\mathrm{NOT}\ \left\{\left(\overline{X}_{j}-\underline{X}_{j}\right)^{{}^{\prime}}\beta_{0}\leq 0\ \mathrm{and}\ \left(\overline{X}_{k}-\underline{X}_{k}\right)^{{}^{\prime}}\beta_{0}\geq 0\ \mathrm{for\ all\ }k\neq j\right\},$
@@
does not imply $\gamma_{j,t,s}\left(\overline{{\bf X}},\underline{{\bf X}}\right)\geq 0$ in the converse direction. This breaks the “if-and-only-if” relationship that the maximum-score criterion function in Manski (1985, 1987) is built upon, and thus the maximum-score estimator does not generalize to multi-index settings. The lack of “if-and-only-if” relationship in the multi-index setting equivalence leads to a key difference in the criterion functions, and consequently a different approach of estimation. Importantly, while the original maximum score criterion and estimator cannot be generalized to multi-index settings, our proposed procedure can be easily applied under a general econometric framework characterized by multi-index single-crossing conditions, which we introduce in Section 4.
@@ 3.1 Two-Step Semiparametric Estimation
We construct our estimator as a semiparametric two-step M-estimator based on (10). The first stage of our procedure concerns with nonparametrically estimating the intertemporal differences in conditional choice probabilities of the following form $\gamma_{j,t,s}\left(\overline{{\bf X}},\underline{{\bf X}}\right)=\mathbb{E}\left[\left.y_{ijt}-y_{ijs}\right|{\bf X}_{i,ts}=\left(\overline{{\bf X}},\underline{{\bf X}}\right)\right]$ for all on-support realizations $\left(\overline{{\bf X}},\underline{{\bf X}}\right)$, all pairs of periods $\left(t,s\right)$ and all products $j$ [^6] In practice, we only need to estimate $\gamma_{j,t,s}$ for $\left(J-1\right)$ products and $\frac{1}{2}T\left(T-1\right)$ ordered pairs of periods. The former is because conditional choice probabilities must sum to one across all $J$ products, so we may easily compute the estimator for the last product from the other $\left(J-1\right)$ estimates: $\gamma_{J,t,s}=1-\sum_{j=1}^{J-1}\gamma_{j,t,s}$. The latter is because $\gamma_{j,t,s}=-\gamma_{j,s,t}$ by construction, so we may estimate it for either $\left(t,s\right)$ or $\left(s,t\right)$. Notice, however, that each ordered pair $\left(t,s\right)$ or $\left(s,t\right)$ provides complementary identifying information, as $\lambda\left({\bf X}_{i,ts};\beta\right)$ and $\lambda\left({\bf X}_{i,st};\beta\right)$ do not admit such kind of deterministic relationship. We note that the first stage estimation includes the observable characteristics of all products $J$. For example, when $J=3$ and $D=3$, there are $3\times 3\times 2=18$ variables in the conditioned set of $\gamma$. Given the potentially large number of regressors, one may want to use neural networks (Bach, 2017; Chen and White, 1999) or penalized sieves (Chen, 2013) for the first step estimation.
@@
$\hat{Q}\left(\beta\right):=\sum_{j=1}^{J}\sum_{t\neq s}^{T}\hat{Q}_{j,t,s}\left(\beta\right),$ $\quad\hat{Q}_{j,t,s}\left(\beta\right):=\frac{1}{N}\sum_{i=1}^{N}G\left(\hat{\gamma}_{j,t,s}\left({\bf X}_{i,ts}\right)\right)\lambda_{j}\left({\bf X}_{i,ts};\beta\right).$
Observing that the scale of $\beta_{0}$ cannot be identified given that $\lambda_{j}\left({\bf X}_{i,ts};\beta\right)$ consists of indicator functions of the the form $\mathbf{\mathbbm{1}}\left\{\left(X_{ijt}-X_{ijs}\right)^{{}^{\prime}}\beta\geq 0\right\}$, we imposes the following scale normalization $\beta_{0}\in\mathbb{\mathbb{S}}^{D-1}:=\left\{v\in\mathbb{R}^{D}:\left\|v\right\|=1\right\}$. Following Chernozhukov, Hong, and Tamer (2007), we define the set estimator by
@@
with $\hat{c}:=O_{p}\left(c_{N}\log N\right)$. We now introduce assumptions for the consistency of $\hat{B}_{\hat{c}}$.
@@
    $\gamma_{j,t,s}\in\Gamma$, and $\mathbb{P}\left(\hat{\gamma}_{j,t,s}\in\Gamma\right)\to 1$, with $\Gamma$ being a $\mathbb{P}$-Donsker class of functions in $L_{2}\left({\bf X}\right)$ such that $\sup_{\gamma_{j,t,s}\in\Gamma}\mathbb{E}\left|\gamma_{j,t,s}\right|<\infty$;
@@
Through Assumption 4 we take as given the large set of theoretical results on nonparametric regression in the literature. Many kernel-based and sieve-based methods have been developed with different properties demonstrated under various sets of conditions. See Wasserman (2006) and Chen (2007) for more comprehensive surveys.
@@
Assumption 5 is not necessary for consistency per se given that our identification result is valid with any choice of the one-sided sign-preserving function $G$, nevertheless we take $G$ to be Lipschitz so as to simplify the proof.
To state the next assumption, we decompose each row (corresponding to each product) of $\overline{{\bf X}}-\underline{{\bf X}}$ as the product of its norm and its direction, i.e., $\overline{X}_{j}-\underline{X}_{j}\equiv r_{j}\left(\overline{{\bf X}}-\underline{{\bf X}}\right)\cdot v_{j}\left(\overline{{\bf X}}-\underline{{\bf X}}\right)$, where $r_{j}\left(\overline{{\bf X}}-\underline{{\bf X}}\right):=\left\|\overline{X}_{j}-\underline{X}_{j}\right\|$, and $v_{j}\left(\overline{{\bf X}}-\underline{{\bf X}}\right):=\left(\overline{X}_{j}-\underline{X}_{j}\right)/\left\|\overline{X}_{j}-\underline{X}_{j}\right\|$ if $\overline{X}_{j}\neq\underline{X}_{j}$ while $v_{j}\left(\overline{{\bf X}}-\underline{{\bf X}}\right):={\bf 0}$ if $\overline{X}_{j}=\underline{X}_{j}$.
@@
The marginal distribution of $v_{j}\left({\bf X}_{it}-{\bf X}_{is}\right)$ has no mass point except possibly at ${\bf 0}$ for each $\left(j,t,s\right)$.
Assumption 6 is a technical assumption that ensures the continuity of the population criterion function. We note that Assumption 6 is fairly weak: it essentially requires that the directions of intertemporal differences in observable characteristics are continuously distributed on their own supports. In particular, this allows all but one dimensions of observable characteristics to be discrete.
With the above assumptions, we now establish the consistency of the set estimator $\hat{B}_{\hat{c}}$ based on Chernozhukov, Hong, and Tamer (2007).
@@
Under Assumptions 26, the set estimator $\hat{B}_{\hat{c}}$ is consistent in Hausdorff distance: $d_{H}\left(\hat{B}_{\hat{c}},B_{0}\right)=o_{p}\left(1\right)$, where $d_{H}\left(\hat{B}_{\hat{c}},B_{0}\right):=\max\left\{\sup_{\beta\in\hat{B}_{\hat{c}}}\inf_{\tilde{\beta}\in B_{0}}\left\|\beta-\tilde{\beta}\right\|,\sup_{\beta\in B_{0}}\inf_{\tilde{\beta}\in\hat{B}_{\hat{c}}}\left\|\beta-\tilde{\beta}\right\|\right\}.$ Furthermore, if $\beta_{0}$ is point-identified on $\mathbb{\mathbb{S}}^{D-1}$, $\left\|\hat{\beta}-\beta_{0}\right\|=o_{p}\left(1\right)$ for any $\hat{\beta}\in\hat{B}:=\arg\min_{\tilde{\beta}\in\mathbb{\mathbb{S}}^{D-1}}\hat{Q}\left(\tilde{\beta}\right)$.

### v3

@@ 2.2 Key Identification Strategy
Relative to the well-known maximum-score criterion function studied by Manski (1985, 1987) under univariate monotonicity, our criterion function is non-standard. This non-standardness arises from a key distinction between multivariate and univariate monotonicity. To see this more clearly, consider the special case of a single-index setting (J=1)(J=1) [^10] This arises naturally in binomial choice models with the characteristics of the outside option set to be zero. In this case, even though there are nominally two choice alternatives, choice behavior is completely determined by a single index based on the characteristics of the non-default option., in which case the following equivalence relationship holds given the univariate monotonicity in the index:
$\left\{\gamma\left({\bf x}_{ts}\right)>0\right\}\ \Leftrightarrow\ \left\{\left(x_{t}-x_{s}\right)^{{}^{\prime}}\beta>0\right\},$ (11)
@@
Such an “if-and-only-if” relationship is a unique feature of the single-index setting that cannot be generalized to the multi-index setting with $J\geq 2$, as the right-hand side of (10),
@@
$\mathrm{NOT}\ \left\{\left(x_{jt}-x_{js}\right)^{{}^{\prime}}\beta_{0}\leq 0,\ \forall j\in\tilde{{\cal J}}\ \mathrm{and}\ \left(x_{kt}-x_{ks}\right)^{{}^{\prime}}\beta_{0}\geq 0\ \forall k\notin\tilde{{\cal J}}\right\}$,
@@
does not imply $\gamma_{\tilde{{\cal J}},ts}\left({\bf x}_{ts}\right)\geq 0$ in the converse direction. This breaks the “if-and-only-if” relationship that the maximum-score criterion function in Manski (1985, 1987) is built upon. Thus, the maximum-score estimator does not generalize to multi-index settings. The lack of “if-and-only-if” relationship in the multi-index setting leads to a key difference in the criterion functions, and consequently a different estimation approach. Importantly, while the original maximum score criterion and estimator cannot be generalized to multi-index settings, our procedure can be applied under a general econometric framework characterized by multi-index single-crossing conditions, which we introduce in Section 4.
2.3 Two-Period Sharpness
We now establish the sharpness of the identified set $B_{0}$ in a two-period setting. This sharpness result can be interpreted as the pairwise sharpness of the identifying restrictions in (10): for fixed periods $s$ and $t$ such that $s<t$, the inequality restrictions in (10) for $(s,t)$ and $(t,s)$ exhaust all the identifying information available from the model (its specification and assumptions) and from the distribution of the observable data in periods $s$ and $t$.
Theorem 2 (Pairwise Sharpness).
Under model (1) and Assumptions 13, $B_{0}$ is sharp for $T=2$.
The proof of Theorem 2 exploits and generalizes a corresponding result in Pakes and Porter (2024). Specifically, Pakes and Porter (2024) consider a specification where the utility index $X_{ijt}^{{}^{\prime}}\beta_{0}$ and an unobserved heterogeneity index $\lambda\left(A_{ijt},\epsilon_{ijt}\right)$ are additively separable, and establish the sharpness of their identification result by showing the existence of a nonnegative solution to a system of linear equations. Here, we consider a more general setup without requiring additive separability and propose a correspondingly more general identification argument. Nevertheless, we show that the sharpness of our identification result under our more general setup can be reduced to the nonnegative solvability of the same system of linear equations in Pakes and Porter (2024). Hence, by the result in Pakes and Porter (2024), our identification result is sharp.
Admittedly, Theorem 2 only establishes sharpness in a two-period setting; however, it does not directly imply “all-period” sharpness for $T\geq 3$. While the existence of an observationally equivalent latent error distribution can be established for any realization ${\bf X}_{i,ts}$ and any pair of periods $\left(t,s\right)$ as in the proof of Theorem 2, to establish the stronger “all-period” sharpness result, we would need to show in addition that there exists an all-period joint distribution of latent errors that matches all-period observed joint conditional choice probabilities and satisfies the pairwise time homogeneity assumption. This appears to be a technically cumbersome exercise, given that the pairwise time-homogeneity assumption enters as an implicit aggregate restriction on the two-period error distributions (with all other periods aggregated out). We thus do not pursue “all-period” sharpness here, and only present the sharpness result above as in Pakes and Porter (2024), which also focuses on two-period (pairwise) sharpness.
@@ 3 Estimation and Computation
3.1 Formulation of Population Criterion Function
We now propose a population criterion function that encodes the identifying information in Proposition 1. We represent the right-hand side of (10) in Boolean algebra by
$\lambda_{\tilde{{\cal J}}}\left({\bf x}_{ts};\beta\right)$ $:=\prod_{k=1}^{J}\mathbf{\mathbbm{1}}\left\{\left(-1\right)^{\mathbf{\mathbbm{1}}\left\{k\in\tilde{{\cal J}}\right\}}\left(x_{kt}-x_{ks}\right)^{{}^{\prime}}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\beta}\geq 0\right\},$ (12)
where $\left(-1\right)^{\mathbf{\mathbbm{1}}\left\{k\in\tilde{{\cal J}}\right\}}$ takes the value 1-1 for $k\in\tilde{{\cal J}}$ and $1$ for $k\notin\tilde{{\cal J}}$. Therefore, Proposition 1 can be written algebraically as: $\gamma_{\tilde{{\cal J}},t,s}\left({\bf x}_{ts}\right)>0$ implies $\lambda_{\tilde{{\cal J}}}\left({\bf x}_{ts};\beta_{0}\right)=0$ for any ${\bf x}_{ts}\in\text{Supp}\left({\bf X}_{i,ts}\right)$.
We now define the following criterion function by taking a cross-sectional expectation over the random realization of ${\bf X}_{i,ts}$ and aggregating over all subsets $\tilde{{\cal J}}\subseteq{\cal J}$:
$Q_{t,s}\left(\beta\right)$ $:=\sum_{\tilde{{\cal J}}\subseteq{\cal J}}\mathbb{E}\left[\mathbf{\mathbbm{1}}\left\{\gamma_{\tilde{{\cal J}},t,s}\left({\bf X}_{i,ts}\right)>0\right\}\lambda_{\tilde{{\cal J}}}\left({\bf X}_{i,ts};\beta\right)\right],$ (13)
which is nonnegative and minimized to zero at $\beta_{0}$. Without normalization and further assumptions for point identification, there could be multiple values of $\beta$ that minimize $Q_{t,s}$ to zero.
More generally, fix any function $G:\mathbb{R}\to\mathbb{R}$ that is one-sided sign preserving, i.e., $G\left(z\right)>0$ for $z>0$ and $G\left(z\right)=0$ for $z\leq 0$. For example, we can choose $G\left(z\right)=\left[z\right]_{+}$ where $\left[z\right]_{+}$ is the positive part function. Then, we define $Q_{t,s}^{G}$ as
$Q_{t,s}^{G}\left(\beta\right)$ $:=\sum_{\tilde{{\cal J}}\subseteq{\cal J}}\mathbb{E}\left[G\left(\gamma_{\tilde{{\cal J}},t,s}\left({\bf X}_{i,ts}\right)\right)\lambda_{\tilde{{\cal J}}}\left({\bf X}_{i,ts};\beta\right)\right],$ (14)
which is also minimized to zero at $\beta_{0}$. The sign-preserving function $G$, if further set to be monotone, continuous, or bounded, serves as a smoothing function that can improve the finite-sample performance of our estimators. We provide more discussions on function $G$ in the next section, when we construct estimators based on the sample analog of the population criterion function defined here.
$Q_{t,s}^{G}$ above is defined for a fixed pair of periods $\left(t,s\right)$, but in practice we may utilize the information across all pairs of periods by defining the aggregated criterion function:
$Q^{G}\left(\beta\right):=\sum_{t\neq s}^{T}Q_{t,s}^{G}\left(\beta\right),\quad\text{for any }\beta\in\mathbb{R}^{D}.$ (15)
For notational simplicity, we suppress $G$ in $Q_{t,s}^{G}$ and $Q^{G}$ in the rest of this paper.
@@ 3.2 Two-Step Semiparametric Estimation
We construct our estimator as a semiparametric two-step M-estimator based on (15). The first stage of our procedure is concerned with nonparametrically estimating the intertemporal differences in conditional choice probabilities of the following form:
$\gamma_{\tilde{{\cal J}},t,s}\left({\bf x}_{ts}\right)=\sum_{j\in\tilde{{\cal J}}}\gamma_{j,t,s}\left({\bf x}_{ts}\right),$
where $\gamma_{j,t,s}\left({\bf x}_{ts}\right)=\mathbb{E}\left[\left.y_{ijt}-y_{ijs}\right|{\bf X}_{i,ts}={\bf x}_{ts}\right]$ can be separately estimated for each product $j\in{\cal J}.$ [^11] In practice, we only need to estimate $\gamma_{j,t,s}$ for $\left(J-1\right)$ products and $\frac{1}{2}T\left(T-1\right)$ ordered pairs of periods. The former is because conditional choice probabilities must sum to one across all $J$ products. Hence, the estimator for the last product from the other $\left(J-1\right)$ estimates can be directly derived by $\gamma_{J,t,s}=-\sum_{j=1}^{J-1}\gamma_{j,t,s}$. The latter is because $\gamma_{j,t,s}=-\gamma_{j,s,t}$ by construction, so we may estimate it for either $\left(t,s\right)$ or $\left(s,t\right)$ pair. Notice, however, that each ordered pair $\left(t,s\right)$ or $\left(s,t\right)$ provides complementary identifying information, as $\lambda\left({\bf X}_{i,ts};\beta\right)$ and $\lambda\left({\bf X}_{i,st};\beta\right)$ do not admit such kind of deterministic relationships. We note that the first stage estimation includes the observable characteristics of all products $J$. For example, when $J=3$ and $D=3$, there are $3\times 3\times 2=18$ variables in the conditioned set of $\gamma$. Given the potentially large number of regressors, one may want to use neural networks (Bach, 2017; Chen and White, 1999) or penalized sieves (Chen, 2013) for the first-step estimation.
@@
$\hat{Q}\left(\beta\right):=\sum_{t\neq s}^{T}\hat{Q}_{\tilde{{\cal J}},t,s}\left(\beta\right),$ $\text{ where }\hat{Q}_{t,s}\left(\beta\right):=\frac{1}{N}\sum_{i=1}^{N}\sum_{\tilde{{\cal J}}\subseteq{\cal J}}G\left(\hat{\gamma}_{\tilde{{\cal J}},t,s}\left({\bf X}_{i,ts}\right)\right)\lambda_{\tilde{{\cal J}}}\left({\bf X}_{i,ts};\beta\right).$
It is worth noting that while $\hat{Q}_{t,s}(\beta)$ is defined as a summation over all $2^{J}$ possible subsets $\tilde{{\cal J}}\subseteq{\cal J}$, computationally there is no need to fully evaluate $Q_{\tilde{{\cal J}},t,s}\left(\beta\right)$ for each possible $\tilde{{\cal J}}\subseteq{\cal J}$ under a given $\beta$ when the knife-edge cases of $(X_{ijt}-X_{ijs})^{{}^{\prime}}\beta=0$ are ignorable. [^12] There are at least two reasons why the “knife-edge” events of the form $(X_{ijt}-X_{ijs})^{{}^{\prime}}\beta=0$ should be ignored. First, $(X_{ijt}-X_{ijs})^{{}^{\prime}}\beta=0$ technically occurs with probability zero for all $\beta$ provided that $X_{ijt}\neq X_{ijs}$ almost surely, which is a natural assumption given that we require $X_{ijt}$ be time-varying. Second, for programming reasons, it is often practically necessary to ignore knife-edge strict equalities of continuously valued variables, since such equalities are extremely sensitive to unavoidable numerical errors induced by the “machine epsilon.” This is because, as long as $\lambda_{\tilde{{\cal J}}}\left({\bf X}_{i,ts};\beta\right)=0$, the contribution from the $\tilde{{\cal J}}-$summand would be zero. However, a careful inspection of $\lambda_{\tilde{{\cal J}}}$ reveals that $\lambda_{\tilde{{\cal J}}}\left({\bf X}_{i,ts};\beta\right)=1$ only if $\tilde{{\cal J}}=\left\{j\in{\cal J}:\left(X_{ijt}-X_{ijs}\right)^{{}^{\prime}}\beta\leq 0\right\}$. Hence, in practical implementation we may simply compute
$\hat{Q}_{t,s}\left(\beta\right):=\frac{1}{N}\sum_{i=1}^{N}G\left(\sum_{j\in{\cal J}}\hat{\gamma}_{j,t,s}\left({\bf X}_{i,ts}\right)\mathbf{\mathbbm{1}}\left\{\left(X_{ijt}-X_{ijs}\right)^{{}^{\prime}}\beta\leq 0\right\}\right).$
In addition, the scale of $\beta_{0}$ is not identified since $\lambda_{j}\left({\bf X}_{i,ts};\beta\right)$ consists of indicator functions of the form $\mathbf{\mathbbm{1}}\left\{\left(X_{ijt}-X_{ijs}\right)^{{}^{\prime}}\beta\geq 0\right\}$. Hence, we impose the scale normalization $\beta_{0}\in\mathbb{\mathbb{S}}^{D-1}:=\left\{v\in\mathbb{R}^{D}:\left\|v\right\|=1\right\}$. Following Chernozhukov, Hong, and Tamer (2007), we define the set estimator by
@@
with $\hat{c}:=O_{p}\left(c_{N}\log N\right)$.
We now introduce assumptions for establishing the consistency of $\hat{B}_{\hat{c}}$.
@@
    $\gamma_{j,t,s}\in\Gamma$, and $\mathbb{P}\left(\hat{\gamma}_{j,t,s}\in\Gamma\right)\to 1$, with $\Gamma$ being a $\mathbb{P}$-Donsker class of functions in $L_{2}\left({\bf X}\right)$.
@@
Through Assumption 4 we take as given the large set of theoretical results on nonparametric regression in the literature. Many kernel-based and sieve-based methods have been developed, with their properties demonstrated under various sets of conditions. See Wasserman (2006) and Chen (2007) for more comprehensive surveys.
@@
Assumption 5 is stronger than necessary for consistency per se given that our identification result is valid with any choice of the one-sided sign-preserving function $G$, nevertheless we take $G$ to be Lipschitz to simplify the proof.
To state the next assumption, we decompose each row (corresponding to each product) of ${\bf x}_{t}-{\bf x}_{s}$ as the product of its norm and its direction, i.e., ${\bf x}_{jt}-{\bf x}_{js}\equiv r_{j}\left({\bf x}_{t}-{\bf x}_{s}\right)v_{j}\left({\bf x}_{t}-{\bf x}_{s}\right)$, where $r_{j}\left({\bf x}_{t}-{\bf x}_{s}\right):=\left\|{\bf x}_{jt}-{\bf x}_{js}\right\|$, and $v_{j}\left({\bf x}_{jt}-{\bf x}_{js}\right):=\left({\bf x}_{jt}-{\bf x}_{js}\right)/\left\|{\bf x}_{jt}-{\bf x}_{js}\right\|$ if ${\bf x}_{jt}\neq{\bf x}_{js}$ while $v_{j}\left({\bf x}_{jt}-{\bf x}_{js}\right):={\bf 0}$ if ${\bf x}_{jt}={\bf x}_{js}$.
@@
The marginal distribution of $v_{j}\left({\bf X}_{it}-{\bf X}_{is}\right)$ has no mass point except possibly at $\bf 0$} and is not supported on any proper linear subspace of $\mathbb{R}^{D}$ for each $\left(j,t,s\right)$ tuple.
Assumption 6 ensures the continuity of the population criterion function. We note that Assumption 6 is mild: it essentially requires that the directions of intertemporal differences in observable characteristics are continuously distributed on their own supports. In particular, this allows all but one dimensions of observable characteristics to be discrete.
With the above assumptions imposed, we now establish the consistency of our set estimator $\hat{B}_{\hat{c}}$, using the results in Chernozhukov, Hong, and Tamer (2007).
@@
Under Assumptions 16, the set estimator $\hat{B}_{\hat{c}}$ is consistent in Hausdorff distance: $d_{H}\left(\hat{B}_{\hat{c}},B_{0}\right)=o_{p}\left(1\right)$, where $d_{H}\left(\hat{B}_{\hat{c}},B_{0}\right)=\max\left\{\sup_{\beta\in\hat{B}_{\hat{c}}}\inf_{\tilde{\beta}\in B_{0}}\left\|\beta-\tilde{\beta}\right\|,\sup_{\beta\in B_{0}}\inf_{\tilde{\beta}\in\hat{B}_{\hat{c}}}\left\|\beta-\tilde{\beta}\right\|\right\}$. Furthermore, if $\beta_{0}$ is point-identified on $\mathbb{\mathbb{S}}^{D-1}$, $\left\|\hat{\beta}-\beta_{0}\right\|=o_{p}\left(1\right)$ for any $\hat{\beta}\in\hat{B}_{\widehat{c}=0}$.